OpenStudy (anonymous):

Write as the logarithm of a single quantity:1/3 log 8 – 2 log 4 + log 6 - 2 log 4 + log 6

jhonyy9 (jhonyy9):

1/3log8=log(2^3)^(1/3) =log2 2log4=log(2^2)^2=log2^4 log6=log6 2log4=log2^4 log6=log6 so your exercise after these will be log2 - log2^4 + log6 - log2^4 + log6 = log(2/(2^4)) +2log6 -log2^4 = =log(1/2^3) +log(36/2^4) = log(1/2^3) +log(36/16) =log(1/8) + log(9/4) = = log(1/8 *9/4) =log(9/32)